# Catalog of outbursts of neutron star LMXBs

### v2

@@ 2 Catalog
We present the catalog of known outbursts of neutron star LMXBs in Table 1, for which we provide a sample here. The full table is in (Table A) in Appendix A). We include a line for each recorded outburst, with a literature reference for each outburst if available. Several objects with repeated bright outbursts can be easily spotted in the public lightcurves of the RXTE PCA Bulge Scan (Swank & Markwardt, 2001)333https://asd.gsfc.nasa.gov/Craig.Markwardt//galscan/main.html, the RXTE All-Sky Monitor (Levine et al., 1996a)444https://heasarc.gsfc.nasa.gov/docs/xte/asm_products.html and/or in MAXI GSC observations (Matsuoka et al., 2009; Suguzaki et al., 2011)555http://maxi.riken.jp/top/index.html. We have added some obvious outbursts visible in these surveys to the list (e.g. outbursts from XTE J1709-267 in 2005, 2009, 2018/19, and 2022). The line for defining an ‘outburst’ can be unclear; for simplicity we require an increase of a factor of 10 from the quiescent flux (if known), and erg/s.
We separate the objects with many outbursts into separate tables, placed in Appendix B. For the most frequently repeating objects with bright outbursts (of order once per year or more), 4U 1608-52, MXB 1730-33 (the Rapid Burster in Liller 1), Aquila X-1, GRS 1747-312 (the eclipser in Terzan 6), and Cir X-1, we use the three public lightcurves above, as well as Swift/XRT public lightcurves (Evans et al., 2007)666https://www.swift.ac.uk/user_objects/, and the INTEGRAL JEM-X Galactic Bulge Monitoring lightcurves (Kuulkers et al., 2007a)777http://integral.esac.esa.int/BULGE/ to develop the best possible list of outbursts from 1996 to 2023, along with literature identification of outbursts before 1996. For the objects with the most frequent outbursts, we also record the MJD date of the outburst peak (as best we can tell). We place the outbursts of NGC 6440 X-2 in a separate table in Appendix B as well, as their high frequency requires MJD dates to differentiate them.
@@ 3.1 Monitoring instruments
RXTE/ASM monitoring produced a dramatic increase in detected outbursts, from January 1996 until 2009-2011 (Levine et al., 1996a)888The RXTE/ASM sensitivity decayed from 2009 until RXTE was turned off in 2011.. RXTE/ASM achieved roughly 15 mCrab ( ergs cm-2 s-1, 3-sigma) sensitivity per day, in 2-10 keV, for known sources. The RXTE/ASM team could identify new sources as faint as 50 mCrab with one day’s data, while integrating for a week could delve down to 7 mCrab, in uncrowded regions (Bradt et al., 2001). Indeed, detections from 1996 on did reach , though the completeness limit was higher, perhaps ergs cm-2 s-1, or 50 mCrab. The RXTE/ASM covered most sources daily, except for about 30 days each year when the Sun was nearest for sources in the ecliptic.
@@
The INTEGRAL satellite (launched 2002, decommissioned in Feb. 2025) has two relevant coded-aperture instruments: IBIS (Ubertini et al., 2003) has peak sensitivity in the 20-60 keV range with a 20 degree field of view (at half sensitivity) and 12’ angular resolution, while JEM-X (Lund et al., 2003) observes in 3-35 keV, with a 7.5 degree field of view and 3.4’ angular resolution. The INTEGRAL Galactic Bulge hard X-ray monitoring program has been particularly productive, with 5-15 mCrab sensitivity per 1800 s observation (every three days), and summed detections reaching down to 2.4 mCrab (Kuulkers et al., 2007a). INTEGRAL could observe the Bulge only 1/3 of the year, for 2 months in each spring and fall. Thirteen of our transients have principal names starting with IGR (only XTE matches this number), suggesting INTEGRAL’s impact. There appears (Fig. 2) to be a small decline in detections in the past few years, which may be due to declining sensitivity of INTEGRAL’s IBIS/ISGRI instrument.999The Integral Galactic Bulge Monitoring Program’s website, http://integral.esac.esa.int/BULGE/, presents 18-40 keV IBIS/ISGRI lightcurves for sources, which show increasing noise from 2016 and stop in 2018.
@@ 3.2 General trends
We can also look at the distribution of peak in detected outbursts (Fig. 4). (We must keep in mind that the majority of these values are likely overestimates, due to the assumed distances being often upper limits, if the observed bursts were not actually Eddington-limited as assumed.) The distribution looks generally similar to the distribution of fluxes, due to the concentration of transient LMXBs in the Galactic Bulge, but due to the variation in distance, the peak is less sharp. An equivalent plot for BH LMXB systems (Fig. 15 in Tetarenko et al. 2016b101010Including a shift down by roughly a factor 3 to move from bolometric to 2-10 keV luminosity.) peaks close to erg/s, with a tail extending down to erg/s. The tendency for BHs to have higher may be attributed to BHs’ bias towards longer orbital periods, which are correlated with higher peak (Wu et al., 2010). At very short periods, even the expected peak luminosities leave black holes in radiatively inefficient states, leading to a strong selection effect against low-luminosity black holes (e.g. Arur & Maccarone 2018). Another selection effect is that verifying a black hole in an LMXB typically requires phase-resolved spectroscopy, which is more difficult in a short-period, thus fainter, system; while verifying a neutron star uses X-ray bursts or pulsations, which aren’t tightly correlated with peak .
@@ Appendix A Outburst details
NGC 6440 X-2 showed numerous weak, short X-ray outbursts in 2009 and 2010, often with a 1-month recurrence time (Heinke et al., 2010)111111Note that Heinke et al. (2010) quote cm-2 in their Table 2, but actually used cm-2.. The RXTE PCA bulge scan reveals at least 7 faint outbursts during 2009-2011, but does not show other similar features except one in 2007, and possibly 4 in 1999. It is not certain whether NGC 6440 X-2 only produced outbursts in 2007 and 2009-2011, or whether we missed others (due perhaps to changes in the outburst peak flux). We list NGC 6440 X-2’s outbursts in a separate table (Table 2), to quote their MJD dates.
@@ Appendix B Frequently outbursting sources
GRS 1747-312 is known to show regular eclipses (in’t Zand et al., 2000), which makes it possible to verify that an outburst from Terzan 6 is produced by this transient, given an accurate orbital ephemeris. Saji et al. (2016) noted a burst, and an absence of a predicted X-ray eclipse, during a Suzaku observation in Sept. 2009, when GRS 1747-312 was not detected in RXTE/PCA monitoring.121212van den Berg et al. (2024) point out that a spike is visible in RXTE/PCA scans on MJD 55088, two days earlier, but that pulsations in the scan reveal that this spike was produced by the known accreting millisecond pulsar IGR J17511–3057, then in outburst. More recently, van den Berg et al. (2024) have shown conclusive evidence that a second eclipsing, bursting X-ray binary, Terzan 6 X2, is located 0.7” from GRS 1747-312. We label the 2009 detection by Suzaku as ”Ter 6 X2” in Table B.

### v3

@@ 2 Catalog
We present the catalog of known outbursts of neutron star LMXBs in Table 1, for which we provide a sample here. The full table is in (Table A) in Appendix A). 333We also maintain an updated version at https://tinyurl.com/5euncwmh. We include a line for each recorded outburst, with a literature reference for each outburst if available. Several objects with repeated bright outbursts can be easily spotted in the public lightcurves of the RXTE PCA Bulge Scan (Swank & Markwardt, 2001)444https://asd.gsfc.nasa.gov/Craig.Markwardt//galscan/main.html, the RXTE All-Sky Monitor (Levine et al., 1996a)555https://heasarc.gsfc.nasa.gov/docs/xte/asm_products.html and/or in MAXI GSC observations (Matsuoka et al., 2009; Suguzaki et al., 2011)666http://maxi.riken.jp/top/index.html. We have added some obvious outbursts visible in these surveys to the list (e.g. outbursts from XTE J1709-267 in 2005, 2009, 2018/19, and 2022). The line for defining an ‘outburst’ can be unclear; for simplicity we require an increase of a factor of 10 from the quiescent flux (if known), and erg/s.
We separate the objects with many outbursts into separate tables, placed in Appendix B. For the most frequently repeating objects with bright outbursts (of order once per year or more), 4U 1608-52, MXB 1730-33 (the Rapid Burster in Liller 1), Aquila X-1, GRS 1747-312 (the eclipser in Terzan 6), and Cir X-1, we use the three public lightcurves above, as well as Swift/XRT public lightcurves (Evans et al., 2007)777https://www.swift.ac.uk/user_objects/, and the INTEGRAL JEM-X Galactic Bulge Monitoring lightcurves (Kuulkers et al., 2007a)888http://integral.esac.esa.int/BULGE/ to develop the best possible list of outbursts from 1996 to 2023, along with literature identification of outbursts before 1996. For the objects with the most frequent outbursts, we also record the MJD date of the outburst peak (as best we can tell). We place the outbursts of NGC 6440 X-2 in a separate table in Appendix B as well, as their high frequency requires MJD dates to differentiate them.
@@ 3.1 Monitoring instruments
RXTE/ASM monitoring produced a dramatic increase in detected outbursts, from January 1996 until 2009-2011 (Levine et al., 1996a)999The RXTE/ASM sensitivity decayed from 2009 until RXTE was turned off in 2011.. RXTE/ASM achieved roughly 15 mCrab ( ergs cm-2 s-1, 3-sigma) sensitivity per day, in 2-10 keV, for known sources. The RXTE/ASM team could identify new sources as faint as 50 mCrab with one day’s data, while integrating for a week could delve down to 7 mCrab, in uncrowded regions (Bradt et al., 2001). Indeed, detections from 1996 on did reach , though the completeness limit was higher, perhaps ergs cm-2 s-1, or 50 mCrab. The RXTE/ASM covered most sources daily, except for about 30 days each year when the Sun was nearest for sources in the ecliptic.
@@
The INTEGRAL satellite (launched 2002, decommissioned in Feb. 2025) has two relevant coded-aperture instruments: IBIS (Ubertini et al., 2003) has peak sensitivity in the 20-60 keV range with a 20 degree field of view (at half sensitivity) and 12’ angular resolution, while JEM-X (Lund et al., 2003) observes in 3-35 keV, with a 7.5 degree field of view and 3.4’ angular resolution. The INTEGRAL Galactic Bulge hard X-ray monitoring program has been particularly productive, with 5-15 mCrab sensitivity per 1800 s observation (every three days), and summed detections reaching down to 2.4 mCrab (Kuulkers et al., 2007a). INTEGRAL could observe the Bulge only 1/3 of the year, for 2 months in each spring and fall. Thirteen of our transients have principal names starting with IGR (only XTE matches this number), suggesting INTEGRAL’s impact. There appears (Fig. 2) to be a small decline in detections in the past few years, which may be due to declining sensitivity of INTEGRAL’s IBIS/ISGRI instrument.101010The Integral Galactic Bulge Monitoring Program’s website, http://integral.esac.esa.int/BULGE/, presents 18-40 keV IBIS/ISGRI lightcurves for sources, which show increasing noise from 2016 and stop in 2018.
@@ 3.2 General trends
We can also look at the distribution of peak in detected outbursts (Fig. 4). (We must keep in mind that the majority of these values are likely overestimates, due to the assumed distances being often upper limits, if the observed bursts were not actually Eddington-limited as assumed.) The distribution looks generally similar to the distribution of fluxes, due to the concentration of transient LMXBs in the Galactic Bulge, but due to the variation in distance, the peak is less sharp. An equivalent plot for BH LMXB systems (Fig. 15 in Tetarenko et al. 2016b111111Including a shift down by roughly a factor 3 to move from bolometric to 2-10 keV luminosity.) peaks close to erg/s, with a tail extending down to erg/s. The tendency for BHs to have higher may be attributed to BHs’ bias towards longer orbital periods, which are correlated with higher peak (Wu et al., 2010). At very short periods, even the expected peak luminosities leave black holes in radiatively inefficient states, leading to a strong selection effect against low-luminosity black holes (e.g. Arur & Maccarone 2018). Another selection effect is that verifying a black hole in an LMXB typically requires phase-resolved spectroscopy, which is more difficult in a short-period, thus fainter, system; while verifying a neutron star uses X-ray bursts or pulsations, which aren’t tightly correlated with peak .
@@ 4 Conclusions
Acknowledgements
We thank the referee for an exceptionally thoughtful and useful report. Swift/XRT lightcurves supplied by the UK Swift Science Centre at the University of Leicester (Evans et al., 2007). RXTE/ASM results provided by the ASM/RXTE teams at MIT and at the RXTE SOF and GOF at NASA’s GSFC. RXTE/PCA results from NASA Goddard Space Flight Center, https://asd.gsfc.nasa.gov/Craig.Markwardt//galscan/main.html. INTEGRAL/JEM-X lightcurves from Kuulkers et al. (2007a) using the OSA10.1 software, distributed by the ISDC (Courvoisier et al., 2003). MAXI results provided by RIKEN, JAXA and the MAXI team (Matsuoka et al., 2009).
This paper made very extensive use of the Astrophysical Data Service (https://ui.adsabs.harvard.edu), ArXiv (http://arxiv.org), and the Astronomer’s Telegram (http://www.astronomerstelegram.org). CH is supported by NSERC Discovery Grant RGPIN-2023-04264. JZ has been supported by a Mitacs Globalink Research Internship.
This paper employs a list of Chandra datasets, obtained by the Chandra X-ray Observatory, contained in the Chandra Data Collection (CDC) 346: https://doi.org/10.25574/cdc.346 (catalog doi:10.25574/cdc.346).
In Appendix A we give the full table (Table A) of data on neutron star LMXB outbursts, except for the frequent outbursters, which are given in Appendix B. Appendix C tabulates information on bursters without recorded outbursts, and Appendix D tabulates the neutron star X-ray binaries that we consider to be persistent systems (not further discussed in this paper).
@@ Appendix A Outburst details
NGC 6440 X-2 showed numerous weak, short X-ray outbursts in 2009 and 2010, often with a 1-month recurrence time (Heinke et al., 2010)121212Note that Heinke et al. (2010) quote cm-2 in their Table 2, but actually used cm-2.. The RXTE PCA bulge scan reveals at least 7 faint outbursts during 2009-2011, but does not show other similar features except one in 2007, and possibly 4 in 1999. It is not certain whether NGC 6440 X-2 only produced outbursts in 2007 and 2009-2011, or whether we missed others (due perhaps to changes in the outburst peak flux). We list NGC 6440 X-2’s outbursts in a separate table (Table 2), to quote their MJD dates.
@@ Appendix B Frequently outbursting sources
GRS 1747-312 is known to show regular eclipses (in’t Zand et al., 2000), which makes it possible to verify that an outburst from Terzan 6 is produced by this transient, given an accurate orbital ephemeris. Saji et al. (2016) noted a burst, and an absence of a predicted X-ray eclipse, during a Suzaku observation in Sept. 2009, when GRS 1747-312 was not detected in RXTE/PCA monitoring.131313van den Berg et al. (2024) point out that a spike is visible in RXTE/PCA scans on MJD 55088, two days earlier, but that pulsations in the scan reveal that this spike was produced by the known accreting millisecond pulsar IGR J17511–3057, then in outburst. More recently, van den Berg et al. (2024) have shown conclusive evidence that a second eclipsing, bursting X-ray binary, Terzan 6 X2, is located 0.7” from GRS 1747-312. We label the 2009 detection by Suzaku as ”Ter 6 X2” in Table B.